Calculate the value of ΔG for the movement of fructose from the portal vein circulation (where its concentration is 1 mM) to the inside of a liver cell (where its concentration is 0.1 mM) at 37°C. Express your answer in units of kJ mol-1 using 2 significant figures.
Answer #1

The expression for the calculation of the free energy released or consumed when the fructose molecule moves from the membrane is:

Jin AG RTIn [F) out

Where,

ΔG is the free energy change

[F]in is the concentration of the fructose inside the liver cell

[F]out is the concentration of the fructose outside the liver cell

R = 8.314 Jmol-1K-1

The conversion of kJ to J is shown below:

1 kJ = 10-3 J

So, R = 8.314* 10-3 kJmol-1K-1

T = 37 0C

The conversion of T(0C) to T(K) is shown below:

T(K) = T(0C) + 273.15

So, the temperature, T = (37 + 273.15) K = 310.15 K

Given,

[F]in = 0.1 mM

[F]out = 1 mM

So,

0.1 AG (8.314x10 kJmol K1(310.15K )ln =-5.94kJmol

Where negative sign signifies spontaneous process.

Answer upto 2 significant digits= -5.9 kJ/mol